$SPAD/src/input list.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool list.output
)set message test on
)set message auto off
)clear all
--S 1 of 33
[2, 4, 5, 6]
--R
--R
--R   (1)  [2,4,5,6]
--R                                                  Type: List(PositiveInteger)
--E 1
--S 2 of 33
[1]
--R
--R
--R   (2)  [1]
--R                                                  Type: List(PositiveInteger)
--E 2
--S 3 of 33
list(1)
--R
--R
--R   (3)  [1]
--R                                                  Type: List(PositiveInteger)
--E 3
--S 4 of 33
append([1,2,3],[5,6,7])
--R
--R
--R   (4)  [1,2,3,5,6,7]
--R                                                  Type: List(PositiveInteger)
--E 4
--S 5 of 33
cons(10,[9,8,7])
--R

--R
--R   (5)  [10,9,8,7]
--R                                                  Type: List(PositiveInteger)
--E 5
)clear all
--S 6 of 33
empty? [x+1]
--R
--R
--R   (1)  false
--R                                                                Type: Boolean
--E 6
--S 7 of 33
([] = nil)@Boolean
--R
--R
--R   (2)  true
--R                                                                Type: Boolean
--E 7
--S 8 of 33
k := [4,3,7,3,8,5,9,2]
--R
--R
--R   (3)  [4,3,7,3,8,5,9,2]
--R                                                  Type: List(PositiveInteger)
--E 8
--S 9 of 33
first k
--R
--R
--R   (4)  4
--R                                                        Type: PositiveInteger
--E 9
--S 10 of 33
k.first
--R
--R
--R   (5)  4
--R                                                        Type: PositiveInteger
--E 10
--S 11 of 33
k.1
--R

--R
--R   (6)  4
--R                                                        Type: PositiveInteger
--E 11
--S 12 of 33
k(1)
--R
--R
--R   (7)  4
--R                                                        Type: PositiveInteger
--E 12
--S 13 of 33
n := #k
--R
--R
--R   (8)  8
--R                                                        Type: PositiveInteger
--E 13
--S 14 of 33
last k
--R
--R
--R   (9)  2
--R                                                        Type: PositiveInteger
--E 14
--S 15 of 33
k.last
--R
--R
--R   (10)  2
--R                                                        Type: PositiveInteger
--E 15
--S 16 of 33
k.(#k)
--R
--R
--R   (11)  2
--R                                                        Type: PositiveInteger
--E 16
)clear all
--S 17 of 33
k := [4,3,7,3,8,5,9,2]
--R

--R
--R   (1)  [4,3,7,3,8,5,9,2]
--R                                                  Type: List(PositiveInteger)
--E 17
--S 18 of 33
k.1 := 999
--R
--R
--R   (2)  999
--R                                                        Type: PositiveInteger
--E 18
--S 19 of 33
k
--R
--R
--R   (3)  [999,3,7,3,8,5,9,2]
--R                                                  Type: List(PositiveInteger)
--E 19
--S 20 of 33
k := [1,2]
--R
--R
--R   (4)  [1,2]
--R                                                  Type: List(PositiveInteger)
--E 20
--S 21 of 33
m := cons(0,k)
--R
--R
--R   (5)  [0,1,2]
--R                                                          Type: List(Integer)
--E 21
--S 22 of 33
m.2 := 99
--R
--R
--R   (6)  99
--R                                                        Type: PositiveInteger
--E 22
--S 23 of 33
m
--R
--R
--R   (7)  [0,99,2]

--R                                                          Type: List(Integer)
--E 23
--S 24 of 33
k
--R
--R
--R   (8)  [99,2]
--R                                                  Type: List(PositiveInteger)
--E 24
)clear all
--S 25 of 33
k := [1,2,3]
--R
--R
--R   (1)  [1,2,3]
--R                                                  Type: List(PositiveInteger)
--E 25
--S 26 of 33
rest k
--R
--R
--R   (2)  [2,3]
--R                                                  Type: List(PositiveInteger)
--E 26
--S 27 of 33
removeDuplicates [4,3,4,3,5,3,4]
--R
--R
--R   (3)  [4,3,5]
--R                                                  Type: List(PositiveInteger)
--E 27
--S 28 of 33
reverse [1,2,3,4,5,6]
--R
--R
--R   (4)  [6,5,4,3,2,1]
--R                                                  Type: List(PositiveInteger)
--E 28
--S 29 of 33
member?(1/2,[3/4,5/6,1/2])
--R
--R
--R   (5)  true

--R                                                                Type: Boolean
--E 29
--S 30 of 33
member?(1/12,[3/4,5/6,1/2])
--R
--R
--R   (6)  false
--R                                                                Type: Boolean
--E 30
)clear all
--S 31 of 33
[1..3,10,20..23]
--R
--R
--R   (1)  [1..3,10..10,20..23]
--R                                         Type: List(Segment(PositiveInteger))
--E 31
--S 32 of 33
expand [1..3,10,20..23]
--R
--R
--R   (2)  [1,2,3,10,20,21,22,23]
--R                                                          Type: List(Integer)
--E 32
--S 33 of 33
expand [1..]
--R
--R
--R   (3)  [1,2,3,4,5,6,7,8,9,10,...]
--R                                                        Type: Stream(Integer)
--E 33
)spool
)lisp (bye)

———-

References^{[1] nothing}

9